OpenStudy (anonymous):

Help me find the solution to these equations please. And explain. 2x - y = 11 3x + 4y = 0

OpenStudy (darronw):

Multiply your top equation so that the y's are the same in both equations. Then add the equations together.

OpenStudy (anonymous):

@DarronW not allowed to do that. I have to use substitution.

OpenStudy (precal):

ok then for sub you need to decide if you want to sub x or y

OpenStudy (anonymous):

I tried both. Can you help me work through it?

OpenStudy (precal):

2x-y=11 I would use y 2x-y=11 -2x -2x to both sides -y=11-2x then I would divide both sides by -1 y=-11+2x

OpenStudy (precal):

sub this into your second equation to solve for y

OpenStudy (precal):

then take your y value and sub it in either equation to solve for x

OpenStudy (anonymous):

Why would you divide both sides by -1?

OpenStudy (anonymous):

@precal

OpenStudy (precal):

because -y is not the same as y

OpenStudy (anonymous):

And it always needs to be in the form of y?

OpenStudy (precal):

yes + y

OpenStudy (anonymous):

So I did that and got to 11x = 4. Now I don't know what to do. I can't have any fractions/decimals

OpenStudy (precal):

3x + 4y = 0 3x+4(-11+2x)=0

OpenStudy (precal):

you did it like this, correct?

OpenStudy (anonymous):

Oops I should've gotten 11x + -44 = 0 right?

OpenStudy (anonymous):

x=4 ?

OpenStudy (anonymous):

x = 4, y = -3, you're right on track.

OpenStudy (precal):

sorry, I don't like to skip steps

OpenStudy (anonymous):

Okay I found the correct answer then. I combied 8x and 3x precal

OpenStudy (anonymous):

Sorry I'm being vague, just trying to work quickly

OpenStudy (precal):

not a problem, you should show all of your steps. It is easy to make mistakes when you are vague......

OpenStudy (anonymous):

Thank you very much. If I have another problem, I will tag you in it if that's okay.

OpenStudy (precal):

ok, I don't know how long I will be online but If I can help you I will